# Storypark: Leveraging Large Language Models to Enhance Children Story Learning Through Child-AI collaboration Storytelling

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computingccs: Human computer interaction (HCI)$c$s: Interactive systems and toolsccs: User interface programming
@@ 2.3. Using LLMs for child-facing storytelling system
LLMS have the potential to facilitate responsive and participatory conversations. They are not limited by a predetermined dialogue paradigm, nor do they rely on pre-written replies. LLMS are able to generate contextually relevant and coherent responses based on the input they receive(Wei et al., 2022), which allows them to achieve near-real conversations, collaborate effectively with children, and create a cohesive response. Generate naturally rich dialogue(Lee et al., 2023). Therefore, llm is very suitable for the field of education and interaction with children. Many researchers have made efforts to incorporate llm into educational technology(Kasneci et al., 2023). These adaptations witness the application of llm in creating educational content(Dijkstra et al., 2022; Gabajiwala et al., 2022), enhancing student engagement and interaction(Abdelghani et al., 2023; Bao, 2019; Tai and Chen, 2023) and personalized learning experiences(Sailer et al., 2023). For example, LLMS have been used to generate children’s narratives(Alabdulkarim et al., 2021; Guan et al., 2020). In different research trajectories, some scholars have used LLMS to create intelligent learning partners that are able to collaborate with humans(Ji et al., 2023), provide feedback(Jia et al., 2021) and encourage students(El Shazly, 2021; Tai and Chen, 2023). A common application involves employing LLMS as dialogue partners in written or oral form, such as in the context of task-oriented dialogues that provide opportunities for language practice(El Shazly, 2021). Or use llm as a story continuation partner to collaborate with children to create stories(Peng et al., 2023). Overall, we believe that large language model applications, can provide children with unique conversational experiences that complement their interactions with others. Our findings have implications for designers and developers of applications based on large language models designed to support children’s learning. In addition, they highlight areas that require further consideration to ensure that the technology is both supportive and effective for different learning experiences.
@@ 4.1.2. Story Continuation
Role setting for large language models In order to better enable LLM’s to continue writing interactive stories, and ensure the establishment of the overall story interaction framework, we first prompting the model to assume the role of a ”interactive story telling robot”. At the beginning of each interaction process, we set the character prompt ”Your character is a children’s story interactive robot, and needs to interact and continue the story based on children’s feedback, and can convey positive values. Your audience is 5-12 years old.”
Role setting for large language models The next part of the prompt focuses on the story context setting. We provide the model with a narrative scenario: ”The main outline of the story and the plot development is [story setting], and through this story teaching the child ’[key value]’. Please expand and continue this story, in the story, design interactive questions 1 and 2 ”. The square brackets [story setting] and [key value] represent the outline of the story development and the educational key value of the story, respectively, where the elements will be dynamically filled according to the story selection. We also set some high-level rules to ensure that the generated story content is ”simple and appropriate for young children” and does not contain anything unrelated to the story.
Story opening template Settings In order for the model to have coherence in story continuation before and after interactive questions, we need to indicate the model continuation requirements. In the continuation phase, the task of the model is to encourage children’s answers and give positive feedback, and then continue the content according to the story outline combined with children’s feedback. This is done with the following detailed prompts: ”Tell the beginning of the story, then continue the story to the position of question 1 and stop to generate interactive question 1, where you need to stop and wait for the child’s feedback”, ”The child’s feedback is [feedback], generate an expansion and continue the story until the child stops at question 2 and waits for the child’s feedback”, ”The child’s feedback is [feedback], generate the ending plot”, and the square brackets [feedback] represent the child’s feedback content.
Context maintenance in story continuations We observed the LLM’s occasionally forgot the storyline which has already expanded as interactions continued and story outline of the setting. This results in LLM’s expands a completely unrelated new story based on the last answer. Since the OpenAI API interface is stateless, we maintain a dialogue context during the interaction, which saves the LLM’s dialogue with the child and the system prompt setting in the current round of interaction. This ensures the consistency of Storypark in story development.
@@ 4.3.2. Evaluation of Content generation
[Dimensions] Here provides a detailed explanation of these metrics:
@@
[…]
Story content evaluation results
[Result] As the figure6 shows, GPT-4 is very close to human-generated content in terms of intelligibility. The content generated by GPT-4 is superior to that of humans in terms of logical rationality, attractiveness, and preference. ChatGPT-4’s attractiveness and explanatory clarity after linking stories to new plot creation may be due to the fact that the task takes advantage of the large language model’s ability to create logical connections of language elements based on specific instructions, and that the logical links made by the large language model are more syntactically tight. Human-generated story content is close to ChatGPT4 in terms of comprehensibility and relevance. Subsequent discussions with the evaluators indicated that the lack of readability of AI-generated stories mainly stems from the fact that the text generated by GPT-4, while consistent in language structure, sometimes lacks internal logical consistency.This observation is consistent with existing research that highlights inherent limitations(Shi et al., 2023; Lovato and Piper, 2019).
@@ 4.3.3. Evaluation of Questions generation
In the problem generation evaluation module, the plot problem and the reason problem have three similar evaluation dimensions. For plot problems, the three dimensions are: understandability, story relevance, and inspiration. The question of reason has the same first two dimensions, and the third different dimension is inclusiveness. Based on these definitions, participants were asked to choose which of two pairs of plot questions or reason questions performed better. [Dimensions] Here provides a detailed explanation of these metrics.
@@
[Result] As the figure7 shows, The evaluations between the human-generated questions and GPT-4 generated questions were close, with no significant differences. This shows that the GPT-4 model’s ability to design problems in story co-creation is acceptable and close to human level. Questions generated by humans score relatively low on story relevance compared to other dimensions. In order to maintain the continuity of the story, the relevance of the problem to the story needs to be considered. This may be because for human participants with early childhood education experience, the task of designing easy questions tends to associate them with various situations in life, resulting in the question design being too divergent and not sufficiently relevant to the story. The questions generated by GPT-4 were also more enlightening and inclusive than those generated by humans, possibly because human participants sometimes created questions that were related to educational goals but unrelated to plot development. The questions generated by GPT-4 can more meaningfully collect children’s feedback on plot development and further promote the development of the story. When it comes to reasoning, human participants sometimes create questions that are related to everyday issues, but unrelated to the core content of the story itself, and can be answered with a clear ”yes” or ”no”, which also limits the inclusiveness of the question and reduces the space for the child to answer.
@@ 5.1.1. Procedure
During the introductory phase, participants will be required to view the three main screens of StoryPark on a Mac computer (i.e., co-creating the plot, understanding, and migrating the creation).
The research assistant first introduced the participants to the use of the buttons in the interface and informed the children about the correct order to use the dialogue buttons and the sketch buttons. The research assistant then explains the meaning of the four icons that imply the flow of the story, and which stage of story creation each corresponds to each of the four nodes.
@@ 5.3.3. Enjoyment
To measure children’s enjoyment in interacting with artificial intelligence or humans, we adapted a questionnaire consisting of 4 items based on the work of Waytz et al. (2014). These questions are as follows:
@@ 6.2.1. Engagement
[Behavioral observation]
@@
[…]
Results of Children’s Engagement in the System
It has been observed from children’s behavior that children generally show a positive attitude toward co-creation with story parks(Figure9) Among the 20 participants, 16 exhibited happy expressions, actions, and positive tones. However, 4 participants displayed confusion and perplexity, which seemed to correlate with the frequency of storytelling and story creation at home. These 4 participants reported listening to stories at home less than three times a week. Following the experiment, 8 participants actively engaged in discussions with the experimenters regarding Storypark’s functionalities. Some child participants even volunteered suggestions for Storypark, which will be addressed in RQ3. The specific experimental data are shown in the figure9. During the whole process of participating in the experiment, whether the expression and movement are happy or not (16:4); Whether participants interact with the story park in a positive tone (16:4); Whether participants were asked to use certain features of the story park repeatedly (12:8); At the end of the experiment, the participants were asked whether they would like to continue discussing the story park with the experimenter (8:12).
@@
[Quantitative analysis] Among the quantitative indicators, all subjects were able to give answers. Among all the recorded answers, 30.6% were meta-comments reflecting the thinking process, and 29.22% contained substantive statements with explanatory significance. The remaining answers that were either irrelevant to the question, or uncertain, or refused to complete the interactive task accounted for 9.8%, and the short answers of less than five words accounted for 39.2%, the highest proportion, possibly because many children answered the plot questions directly with a single word. Therefore, in terms of the combination of quantitative indicators and behavioral observation, most of the subjects actively participated in the process of co-creating stories and interacting with Storypark.
@@ 6.2.2. Enjoyment
From the figure10, it can be seen that children have an overall positive perception of learning story content through collaborative storytelling. There is a general high level of engagement and satisfaction: children’s willingness and satisfaction in using the AI story assistant were assessed using a four-dimensional scale ranging from 1 to 5. Among the 20 participants, 16 gave a rating of 5 for all four dimensions on the scale. Children expressed high satisfaction with the combination of voice and visual modalities. Two participants, aged 5-6, accurately articulated the value of the AI story assistant without guidance from experimenters: ”It can save my parents’ energy because it is a machine and can tell me new stories unlimited times” and ”My dad is very busy with work, he only tells me once, but AI can tell me many times”.
@@ 7.2.1. Impact of LLM Unpredictability Output
During the development process of interactive storytelling, we discovered that LLMs exhibit a high degree of unpredictability in their output. Slight variations in wording or syntax within the story outlines set in prompts can significantly affect the content generated by the LLMs, with each iteration of story generation producing variations. Modifications in prompt engineering cannot consistently guarantee the desired outcomes, as the interpretation of our story outlines and prompt details by LLMs remains a ”black-box” process of input and output. LLMs exhibit “dispersion” in their potential predictions for a fixed input (i.e., prompt) has also been noted by Lin et al. in their research. To ensure that prompt settings yield relatively stable and consistent outputs, extensive testing of story generation was conducted during development to assure the stability of the story continuations.
@@ 7.2.2. Fine-grained Control of LLMs’ Performance
LLMs demonstrate competence in understanding and executing simple directives conveyed through basic descriptive language. However, it remains a challenge when tasked with precise instructions that incorporate numerical details. Our development experience revealed that LLMs can effectively and consistently execute instructions lacking precise descriptions, such as constraints related to plot continuation or requirements tailored to specific story audience demographics. But when faced with numerically precise instructions, such as ”The story’s opening should exceed 100 words but not surpass 200 words,” LLMs frequently disregard these word count constraints, producing content shorter than 100 words. This indicates that current LLMs still have limited capabilities in understanding numerical values and reasoning, necessitating the integration of validation layers (He-Yueya) to ensure model outputs align with exact requirements. Feedback loops (Lee) also emerges as a viable strategy.
Additionally, we encountered issues with LLMs deviating from the intended story progression, wherein the model would overlook the initial story outline and plot development constraints. To mitigate this, we employed conversational contexts to maintain a history record, and removed dialogue records from the keyword extraction phase to prevent interference with story generation, Shi et al. also found this phenomenon. For longer stories, injection prompting at the end of each input has been suggested (Liu et al.) to enhance LLM performance.
@@ 7.2.3. Impact of LLM Hallucinations
Hallucinations in LLMs refer to instances where the generated responses possess internally flawed logic or absurd content, albeit appearing plausible (Li Ji). At times, LLMs may not comprehend the real-world logic of object interactions within story continuations or whether children’s feedback aligns with the question requirements, merely producing content that seems correct based on the constraints of our prompt engineering. We observed hallucinatory phenomena in our story continuations, for example, in a story about a turtle, children suggested using the turtle shell to make herbal jelly, and the LLM described the process of making herbal jelly from the turtle shell, despite the fact that turtle shells cannot be used to make herbal jelly and such a storyline cannot proceed further.

### v2

@@ Abstract.
$c$s: Human-centered computingccs: Human computer interaction (HCI)$c$s: Interactive systems and toolsccs: User interface programming
@@ 2.3. Using LLMs for child-facing storytelling system
LLMS have the potential to facilitate responsive and participatory conversations. They are not limited by a predetermined dialogue paradigm, nor do they rely on pre-written replies. LLMS are able to generate contextually relevant and coherent responses based on the input they receive(Wei et al., 2022), which allows them to achieve near-real conversations, collaborate effectively with children, and create a cohesive response. Generate naturally rich dialogue(Lee et al., 2023a). Therefore, llm is very suitable for the field of education and interaction with children. Many researchers have made efforts to incorporate llm into educational technology(Kasneci et al., 2023). These adaptations witness the application of llm in creating educational content(Dijkstra et al., 2022; Gabajiwala et al., 2022), enhancing student engagement and interaction(Abdelghani et al., 2023; Bao, 2019; Tai and Chen, 2023) and personalized learning experiences(Sailer et al., 2023). For example, LLMS have been used to generate children’s narratives(Alabdulkarim et al., 2021; Guan et al., 2020). In different research trajectories, some scholars have used LLMS to create intelligent learning partners that are able to collaborate with humans(Ji et al., 2023a), provide feedback(Jia et al., 2021) and encourage students(El Shazly, 2021; Tai and Chen, 2023). A common application involves employing LLMS as dialogue partners in written or oral form, such as in the context of task-oriented dialogues that provide opportunities for language practice(El Shazly, 2021). Or use llm as a story continuation partner to collaborate with children to create stories(Peng et al., 2023). Overall, we believe that large language model applications, can provide children with unique conversational experiences that complement their interactions with others. Our findings have implications for designers and developers of applications based on large language models designed to support children’s learning. In addition, they highlight areas that require further consideration to ensure that the technology is both supportive and effective for different learning experiences.
@@ 4.1.2. Story Continuation
Role setting for large language models: In order to better enable LLM’s to continue writing interactive stories, and ensure the establishment of the overall story interaction framework, we first prompting the model to assume the role of a ”interactive story telling robot”. At the beginning of each interaction process, we set the character prompt ”Your character is a children’s story interactive robot, and needs to interact and continue the story based on children’s feedback, and can convey positive values. Your audience is 5-12 years old.”
Story opening template Settings: The next part of the prompt focuses on the story context setting. We provide the model with a narrative scenario: ”The main outline of the story and the plot development is [story setting], and through this story teaching the child ’[key value]’. Please expand and continue this story, in the story, design interactive questions 1 and 2 ”. The square brackets [story setting] and [key value] represent the outline of the story development and the educational key value of the story, respectively, where the elements will be dynamically filled according to the story selection. We also set some high-level rules to ensure that the generated story content is ”simple and appropriate for young children” and does not contain anything unrelated to the story.
Story continuation template Settings: In order for the model to have coherence in story continuation before and after interactive questions, we need to indicate the model continuation requirements. In the continuation phase, the task of the model is to encourage children’s answers and give positive feedback, and then continue the content according to the story outline combined with children’s feedback. This is done with the following detailed prompts: ”Tell the beginning of the story, then continue the story to the position of question 1 and stop to generate interactive question 1, where you need to stop and wait for the child’s feedback”, ”The child’s feedback is [feedback], generate an expansion and continue the story until the child stops at question 2 and waits for the child’s feedback”, ”The child’s feedback is [feedback], generate the ending plot”, and the square brackets [feedback] represent the child’s feedback content.
Context maintenance in story continuations: We observed the LLM’s occasionally forgot the storyline which has already expanded as interactions continued and story outline of the setting. This results in LLM’s expands a completely unrelated new story based on the last answer. Since the OpenAI API interface is stateless, we maintain a dialogue context during the interaction, which saves the LLM’s dialogue with the child and the system prompt setting in the current round of interaction. This ensures the consistency of Storypark in story development.
@@ 4.3.2. Evaluation of Content generation
Dimensions: Here provides a detailed explanation of these metrics:
@@
Result: As the figure6 shows, GPT-4 is very close to human-generated content in terms of intelligibility. The content generated by GPT-4 is superior to that of humans in terms of logical rationality, attractiveness, and preference. ChatGPT-4’s attractiveness and explanatory clarity after linking stories to new plot creation may be due to the fact that the task takes advantage of the large language model’s ability to create logical connections of language elements based on specific instructions, and that the logical links made by the large language model are more syntactically tight. Human-generated story content is close to ChatGPT4 in terms of comprehensibility and relevance. Subsequent discussions with the evaluators indicated that the lack of readability of AI-generated stories mainly stems from the fact that the text generated by GPT-4, while consistent in language structure, sometimes lacks internal logical consistency.This observation is consistent with existing research that highlights inherent limitations(Shi et al., 2023; Lovato and Piper, 2019).
@@ 4.3.3. Evaluation of Questions generation
In the problem generation evaluation module, the plot problem and the reason problem have three similar evaluation dimensions. For plot problems, the three dimensions are: understandability, story relevance, and inspiration. The question of reason has the same first two dimensions, and the third different dimension is inclusiveness. Based on these definitions, participants were asked to choose which of two pairs of plot questions or reason questions performed better.
Dimensions: Here provides a detailed explanation of these metrics.
@@
Result: As the figure7 shows, The evaluations between the human-generated questions and GPT-4 generated questions were close, with no significant differences. This shows that the GPT-4 model’s ability to design problems in story co-creation is acceptable and close to human level. Questions generated by humans score relatively low on story relevance compared to other dimensions. In order to maintain the continuity of the story, the relevance of the problem to the story needs to be considered. This may be because for human participants with early childhood education experience, the task of designing easy questions tends to associate them with various situations in life, resulting in the question design being too divergent and not sufficiently relevant to the story. The questions generated by GPT-4 were also more enlightening and inclusive than those generated by humans, possibly because human participants sometimes created questions that were related to educational goals but unrelated to plot development. The questions generated by GPT-4 can more meaningfully collect children’s feedback on plot development and further promote the development of the story. When it comes to reasoning, human participants sometimes create questions that are related to everyday issues, but unrelated to the core content of the story itself, and can be answered with a clear ”yes” or ”no”, which also limits the inclusiveness of the question and reduces the space for the child to answer.
@@ 5.1.1. Procedure
During the introductory phase, participants will be required to view the three main screens of StoryPark on a Mac computer (i.e., co-creating the plot, understanding, and migrating the creation). The research assistant first introduced the participants to the use of the buttons in the interface and informed the children about the correct order to use the dialogue buttons and the sketch buttons. The research assistant then explains the meaning of the four icons that imply the flow of the story, and which stage of story creation each corresponds to each of the four nodes.
@@ 5.3.3. Enjoyment
To measure children’s enjoyment in interacting with artificial intelligence or humans, we adapted a questionnaire consisting of 4 items based on the prior work of Waytz(Waytz et al., 2014). These questions are as follows:
@@ 6.2.1. Engagement
Behavioral observation: It has been observed from children’s behavior that children generally show a positive attitude toward co-creation with story parks(Figure9)
@@
Among the 20 participants, 16 exhibited happy expressions, actions, and positive tones. However, 4 participants displayed confusion and perplexity, which seemed to correlate with the frequency of storytelling and story creation at home. These 4 participants reported listening to stories at home less than three times a week. Following the experiment, 8 participants actively engaged in discussions with the experimenters regarding Storypark’s functionalities. Some child participants even volunteered suggestions for Storypark, which will be addressed in RQ3. The specific experimental data are shown in the figure9. During the whole process of participating in the experiment, whether the expression and movement are happy or not (16:4); Whether participants interact with the story park in a positive tone (16:4); Whether participants were asked to use certain features of the story park repeatedly (12:8); At the end of the experiment, the participants were asked whether they would like to continue discussing the story park with the experimenter (8:12).
@@
Quantitative analysis: Among the quantitative indicators, all subjects were able to give answers. Among all the recorded answers, 30.6% were meta-comments reflecting the thinking process, and 29.22% contained substantive statements with explanatory significance. The remaining answers that were either irrelevant to the question, or uncertain, or refused to complete the interactive task accounted for 9.8%, and the short answers of less than five words accounted for 39.2%, the highest proportion, possibly because many children answered the plot questions directly with a single word. Therefore, in terms of the combination of quantitative indicators and behavioral observation, most of the subjects actively participated in the process of co-creating stories and interacting with Storypark.
@@ 6.2.2. Enjoyment
It can be seen that children have an overall positive perception of learning story content through collaborative storytelling. There is a general high level of engagement and satisfaction: children’s willingness and satisfaction in using the AI story assistant were assessed using a four-dimensional scale ranging from 1 to 5. Among the 20 participants, 16 gave a rating of 5 for all four dimensions on the scale. Children expressed high satisfaction with the combination of voice and visual modalities. Two participants, aged 5-6, accurately articulated the value of the AI story assistant without guidance from experimenters: ”It can save my parents’ energy because it is a machine and can tell me new stories unlimited times” and ”My dad is very busy with work, he only tells me once, but AI can tell me many times”.
@@ 7.2.1. Impact of LLM Unpredictability Output
During the development process of interactive storytelling, we discovered that LLMs exhibit a high degree of unpredictability in their output. Slight variations in wording or syntax within the story outlines set in prompts can significantly affect the content generated by the LLMs, with each iteration of story generation producing variations. Modifications in prompt engineering cannot consistently guarantee the desired outcomes, as the interpretation of our story outlines and prompt details by LLMs remains a ”black-box” process of input and output. LLMs exhibit “dispersion” in their potential predictions for a fixed input (i.e., prompt) has also been noted by Lin et al. in their research(Lin et al., 2023). To ensure that prompt settings yield relatively stable and consistent outputs, extensive testing of story generation was conducted during development to assure the stability of the story continuations.
@@ 7.2.2. Fine-grained Control of LLMs’ Performance
LLMs demonstrate competence in understanding and executing simple directives conveyed through basic descriptive language. However, it remains a challenge when tasked with precise instructions that incorporate numerical details. Our development experience revealed that LLMs can effectively and consistently execute instructions lacking precise descriptions, such as constraints related to plot continuation or requirements tailored to specific story audience demographics. But when faced with numerically precise instructions, such as ”The story’s opening should exceed 100 words but not surpass 200 words,” LLMs frequently disregard these word count constraints, producing content shorter than 100 words. This indicates that current LLMs still have limited capabilities in understanding numerical values and reasoning, necessitating the integration of validation layers(He-Yueya et al., 2023) to ensure model outputs align with exact requirements. Feedback loops (Lee et al., 2023b) also emerges as a viable strategy.
Additionally, we encountered issues with LLMs deviating from the intended story progression, wherein the model would overlook the initial story outline and plot development constraints. To mitigate this, we employed conversational contexts to maintain a history record, and removed dialogue records from the keyword extraction phase to prevent interference with story generation, Shi also found this phenomenon(Shi et al., 2023). For longer stories, injection prompting at the end of each input has been suggested to enhance LLM performance(Liu et al., 2024).
@@ 7.2.3. Impact of LLM Hallucinations
Hallucinations in LLMs refer to instances where the generated responses possess internally flawed logic or absurd content, albeit appearing plausible(Ji et al., 2023b; Li et al., 2023). At times, LLMs may not comprehend the real-world logic of object interactions within story continuations or whether children’s feedback aligns with the question requirements, merely producing content that seems correct based on the constraints of our prompt engineering. We observed hallucinatory phenomena in our story continuations, for example, in a story about a turtle, children suggested using the turtle shell to make herbal jelly, and the LLM described the process of making herbal jelly from the turtle shell, despite the fact that turtle shells cannot be used to make herbal jelly and such a storyline cannot proceed further.